OpenStudy (anonymous):

Find f(x) and g(x) so that the function can be described as y = f(g(x)). y= 8/x^2 +2

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):



OpenStudy (anonymous):

@jim_thompson5910 last problem!

OpenStudy (anonymous):

@geerky42 please :(

jimthompson5910 (jim_thompson5910):

were you able to get anywhere?

OpenStudy (anonymous):

no idea. the y messed me up

jimthompson5910 (jim_thompson5910):

what would happen if you replaced x^2 with x

jimthompson5910 (jim_thompson5910):

and made it \[\Large \frac{8}{x} + 2\]

OpenStudy (anonymous):

square on the other side?

jimthompson5910 (jim_thompson5910):

what does the notation f(g(x)) mean

OpenStudy (anonymous):

you find g(x) and then plug in for f(x)

OpenStudy (anonymous):

@geerky42

jimthompson5910 (jim_thompson5910):

so let's say we had f(x) = x+2 and g(x) = 10x what is f(g(x))

OpenStudy (anonymous):

10x^2+20x

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

you just plug in g(x) into x of f(x)

jimthompson5910 (jim_thompson5910):